Dolomite, [CaMg(CO3)2] is found in a soil sample. A geochemist titrates 24.65 g of soil with 57.85 mL of 0.3315M HCl. What is the mass % of dolomite in the soil?
Answer #1

Dolomite, [CaMg(CO3)2] is found in a soil sample. A geochemist titrates 24.65 g of soil with 57.85 mL of 0.3315M HCl. What is the mass % of dolomite in the soil?

2H+ + CO3-2 --> H2CO3

mol of HCl = Macid*Vacid = 57.85*0.3315 = 19.177275 mmol of HCl = 19.177275*10^-3 mol

now,

ratio is 2:!

19.177275 mmol of HCl = 19.177275/2 = 9.5886375 mmol of CO3-2

[CaMg(CO3)2]

1 mol of dolomite = 2 mol fo CO3-2

x mol of dolomit = 9.5886375 mmol of CO2-3

x = 9.5886375 /2 = 4.79431 mmol of dolomite

mmol of CaMg(CO3)2 --> 4.79431

MW of dolmite = 184.40 g/mol

mass of dolomite = (4.79431 *10^-3)(184.40 ) =0.884 g of CaMg(CO3)2

then

% mass = 0.884/24.65*100% = 3.5862%